OpenStudy (brinazarski):

Find the remainder when x^13 + 1 is divided by x - 1. Please explain!

OpenStudy (brinazarski):

If it matters, I haven't done this stuff in awhile >.<

Parth (parthkohli):

When \(p(x)\) is divided by \(x - k\), then the remainder is \(p(k)\).

Parth (parthkohli):

Here, just find \(1^{13} + 1\)

OpenStudy (brinazarski):

... 2?

OpenStudy (anonymous):

here u have to use factor and remainder theorem let f(x) =x^13 +1 if it is divided by x-1 then x^13 +1= (x-1)Q(x) +R(x) hence putting both sides x=1 we have R(1) =1^13+1 =2 hence the remainder is 2 when

OpenStudy (anonymous):

as the term (x-1)Q(x) vanishes when x=1

OpenStudy (brinazarski):

I'm confused... not good with math. But the "p(x) is divided by x−k, then the remainder is p(k)." applies to all situations, right? Cuz I'll just jot that down for my own notes....

OpenStudy (anonymous):

yups

OpenStudy (anonymous):

see F(x)=(x-a)Q(x) +R(x) where Q(x) is the quotient and R(x) is the remainder thus if u put x=a then F(xa=(a-a)Q(a) +R(a) hence R(a) =F(a) and hence it applies to all situations

OpenStudy (brinazarski):

Ah... sort of makes sense. Thank you! :)

OpenStudy (anonymous):

welcome